In a maize coloured endosperm (C) is dominant over colourless (c) and full endosperm (R) is dominant over shrunken (r). When a dihybrid of F1 generation was test crossed, it produced four phenotypes in the following %:

1. Coloured and full - 45%

2. coloured and shrunken - 5%

3. Colourless - Full - 4%

4. Colourless - shrunken - 46%

From these data what would be distance between the two non allelic genes?
Answer #1

Representation:

Trait 1- Endosperm color:

  • Colored = C = dominant, genotype= CC/Cc
  • Colorless = c= recessive, genotype= cc

Trait 2- Endosperm Type:

  • Full= R = dominant, genotype= RR/Rr
  • Shrunken = r= recessive, genotype= rr

CCRR x ccrr

F1: CcRr

  • Test cros:

Cross: CcRr x ccrr

For CcRr- gcmetes= CR, cr, Cr, cR

For ccrr-gcmetes = cr

cr

cr

CR

CcRr

CcRr

Cr

Ccrr

Ccrr

cR

ccRr

ccRr

cr

ccrr

ccrr

For independent assortment: ratio= 1:1:1:1

Recombinants =Colored-shrunken =5 % and Colorless-full = 4% =9%

Parental = Colored-full = 45% and Colorless-shrunken =46% =45+46= 91%

Total =100

RF= 9/100 x 100 = 9%

Distance = 9 MU